OpenStudy (anonymous):

find the slope intercept equation of the line that has the given characterstics. slope -7 y-intercept (0,8) the equation is y= ?

OpenStudy (anonymous):

-7

OpenStudy (triciaal):

y = mx + c 8 =-7(0) + c c = 8 + 7 = 15 y = -7 x + 15

OpenStudy (triciaal):

Isaiah250 question said find the slope intercept eqn of the line cannot just be a number

OpenStudy (anonymous):

oh yea didn't read that

OpenStudy (triciaal):

btw you gave the slope, they already gave y = 8 from the point (0,8)

OpenStudy (anonymous):

@Isaiah250

OpenStudy (anonymous):

@triciaal

OpenStudy (triciaal):

to cmenendez55 are you ok now?

OpenStudy (anonymous):

@cmenendez55

OpenStudy (anonymous):

y=-7x+15 is wromg

OpenStudy (anonymous):

what is it?

OpenStudy (anonymous):

i dont know but its wrong my hw is online and it said its wrong

OpenStudy (anonymous):

ohhh

OpenStudy (acxbox22):

the answer is y=-7x+8

OpenStudy (acxbox22):

since the coefficient of x is always the slope in linear equations and b or y-intercept is equal to 8 this is the equation

OpenStudy (anonymous):

omg thank u

OpenStudy (triciaal):

the method is correct and y = mx + c

OpenStudy (acxbox22):

it is y=mx+b

OpenStudy (triciaal):

ok y = mx + b here I forget sometimes my son corrected me before too i learnt c for constant sorry